OpenStudy (anonymous):

in a triangle ABC angle A=40 and angle B=57 and side c is 30 cm find angle C

OpenStudy (anonymous):

i got the angle C its 83

OpenStudy (anonymous):

find the length of a and b

OpenStudy (anonymous):

@ALLTehMaffs

OpenStudy (anonymous):

law of sines \[\frac{a}{\sin A} = \frac{c}{\sin C}\] \[\frac{b}{\sin B} = \frac{c}{\sin C}\]

OpenStudy (anonymous):

a=19.47 b=25.41

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

I got b=25.35 , which is really close so I dunno what's up

OpenStudy (anonymous):

how about a?

OpenStudy (anonymous):

that one I got the same

OpenStudy (anonymous):

er, I got 19.43

OpenStudy (anonymous):

find the area of the triangle?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

i dnt understand

OpenStudy (anonymous):

I did it wrong, 1 sec.

OpenStudy (anonymous):

herons law?

OpenStudy (anonymous):

are you there?

OpenStudy (anonymous):

Area of triangle = 1/2 (Base)(Height) = (1/2) (c)(h)  Make right triangle with b as hypotenuse to find the height, h SORRY \[h=b \cos A\] \[Area=\frac{1}{2}ch\] \[ = \frac{1}{2}c(b \cos A)\]